# Attention-Guided Perturbation for Unsupervised Image Anomaly Detection

### v2

@@ 4.7 Extending to Few-shot Anomaly Detection
To further demonstrate the generality and potential of our framework, we conduct experiments on MVTec-AD under the challenging few-shot anomaly detection setups. We compared our method with some competitive methods and the models designed specifically for few-shot setting, including SPADE Cohen and Hoshen (2020), PaDiM Defard et al. (2021), PatchCore Roth et al. (2022), WinCLIP+ Jeong et al. (2023), RWDA Tamura (2023), FastRcon Fang et al. (2023), PromptAD Li et al. (2024). As shown in Table 10, our method also obtains a leading few-shot anomaly detection performance, especially at pixel level. Specifically, our method achieves the best localization performance of 96.2% and 97.3% P-AUC under 2-shot and 4-shot settings. Compared to the competitive method PromptAD, which depends on a large pretrained CLIP model and intricate prompt tuning, we achieved a 0.8% P-AUC increase under 4-shot settings. Patchcore is a competitive anomaly detection model that relies on a resource-intensive memory bank to store nominal patch features. However, when faced with the more challenging few-shot setup, our approach remarkably outperforms it, with improvements of 6.4%, 7.0%, and 9.0% in I-AUC, and 3.0%, 2.9%, and 3.6% in P-AUC under the 4-shot, 2-shot, and 1-shot settings, respectively. The outstanding performance of our framework in few-shot settings highlights the unique advantages of our attention-guided perturbation strategy for scenarios with severely limited data. This success is rooted in the strategy’s ability to prioritize perturbations in semantically critical regions of the input samples, which fundamentally accelerates model convergence and maximizes the utility of every available training sample. By dynamically focusing on areas that are most informative for modeling normal distribution, the attention-guided perturbation method ensures robust normal distribution modeling even with minimal training data, making it suited for few-shot setting.

### v3

@@ 4.6 Extending to Few-shot Anomaly Detection
To further demonstrate the generality and potential of our framework, we conduct experiments on MVTec-AD under the challenging few-shot anomaly detection setups. We compared our method with some competitive methods and the models designed specifically for few-shot setting, including SPADE [65], PaDiM [2], PatchCore [35], WinCLIP+ [49], RWDA [66], FastRcon [67], PromptAD [51]. As shown in Table 9, our method also obtains a leading few-shot anomaly detection performance, especially at pixel level. Specifically, our method achieves the best localization performance of 96.2% and 97.3% P-AUC under 2-shot and 4-shot settings. Compared to the competitive method PromptAD, which depends on a large pretrained CLIP model and intricate prompt tuning, we achieved a 0.8% P-AUC increase under 4-shot settings. Patchcore is a competitive anomaly detection model that relies on a resource-intensive memory bank to store nominal patch features. However, when faced with the more challenging few-shot setup, our approach remarkably outperforms it, with improvements of 6.4%, 7.0%, and 9.0% in I-AUC, and 3.0%, 2.9%, and 3.6% in P-AUC under the 4-shot, 2-shot, and 1-shot settings, respectively. The outstanding performance of our framework in few-shot settings highlights the unique advantages of our attention-guided perturbation strategy for scenarios with severely limited data.
@@ 6 Conclusion
7 Acknowledge
This work is supported by the National Natural Science Foundation of China (No.62302188); Hubei Province Natural Science Foundation (No.2023AFB267); Fundamental Research Funds for the Central Universities (No.2662023XXQD001).